# From small eigenvalues to large cuts, and Chowla’s cosine problem

### v1

@@ 11 Densification — Phase 4
First, assume that $N_{2}\leq\frac{1}{8}pn^{2}$. Note that $\mu_{1}^{2}+P_{2}+N_{2}=\|B\|_{F}^{2}$ is twice the number of edges of $\overline{G}$, so $\mu_{1}^{2}+P_{2}+N_{2}=2p\binom{n}{2}$, from which
$P_{2}\geq pn^{2}/2-\mu_{1}^{2}-N_{2}\geq pn^{2}/2-C^{2}p^{2}n^{2}-pn^{2}/8\geq pn^{2}/4,$
where we have used that $pC^{2}\leq 10^{-3}$ in the last inequality. But then $\operatorname{surp}^{*}(G)=\Omega(p^{1/2}n^{3/2})$ by the second highlighted inequality, and we are done.
Hence, in the rest of the proof, we may assume that $N_{2}\geq\frac{1}{8}pn^{2}$. By the inequality between the quadratic and cubic mean, we have
@@
Next, consider the quantity $T=N_{3}-P_{3}$. Observe that $\mu_{1}^{3}-T=\sum_{i=1}^{n}\mu_{i}^{3}$ is six-times the number of triangles of $\overline{G}$. In particular, $\mu_{1}^{3}-T$ it is nonnegative, showing that $T\leq\mu_{1}^{3}\leq\bar{\Delta}^{3}$. Assume that $N_{3}\geq 2T$, then $P_{3}\geq N_{3}/2$. By the third highlighted inequality, we then have
$\operatorname{surp}^{*}(G)\geq\Omega\left(\frac{P_{3}}{\bar{\Delta}}\right)\geq\Omega\left(\frac{N_{3}}{\bar{\Delta}}\right)\geq\Omega(C^{-1}p^{1/2}n^{3/2}).$
Hence, we are done in this case as well.
@@
Finally, assume that $N_{3}\leq 2T$, then $\bar{\Delta}^{3}\geq T\geq N_{3}/2$. By the Cauchy-Schwartz inequality applied to the sequences $(|\mu_{i}|^{3})_{\mu_{i}<0}$ and $(|\mu_{i}|)_{\mu_{i}<0}$, we have the inequality $N_{1}N_{3}\geq N_{2}^{2}$, which gives
$N_{1}\geq\frac{N_{2}^{2}}{N_{3}}\geq\frac{p^{2}n^{4}}{128\bar{\Delta}^{3}}\geq\frac{n}{128C^{3}p}.$
But $0=\operatorname{tr}(B)=\mu_{1}+P_{1}-N_{1}$, from which $P_{1}=N_{1}-\mu_{1}\geq\frac{n}{128C^{3}p}-\bar{\Delta}\geq\frac{n}{500C^{3}p}$. Hence, as $\operatorname{surp}^{*}(G)\geq P_{1}$, we are done. ∎
Next, we present a simple technical lemma which shows that every graph contains a large induced $O(\log n)$-balanced graph.
Lemma 11.3.
Let $G$ be an $n$-vertex graph of edge density $p$, and let $C\geq 4\log_{2}n$. Then $G$ contains a $C$-balanced induced subgraph on at least $(1-2\log_{2}n/C)n$ vertices of edge density at most $p$.
@@
Let $d=d(G)$ be the average degree of $G$. We perform an inductive process, where in each step we delete the vertices of high degree. More precisely, let $G_{0}=G$, and define the sequence of induced subgraphs $G_{0}\supset G_{1}\supset$... as follows. Having already defined $G_{i}$, we denote its number of vertices by $n_{i}$ and its average degree by $d_{i}$. To define $G_{i+1}$, remove from $G_{i}$ all vertices of degree at least $Cd_{i}/2$, if any such vertices exist. The process halts once either all vertices of $G_{i}$ have degree less than $Cd_{i}/2$, or the average degree of $G_{i}$ is at least $d_{i-1}/2$. Let $I$ be the last index $i$ for which $G_{i}$ is defined.
For each $i$, it is not hard to verify that the density of $G_{i}$ is smaller than the density of $G_{i-1}$. Moreover, since the average degree of $G_{i}$ is $d_{i}$, there are at most $2n_{i}/C$ vertices of degree larger than $Cd_{i}/2$ in $G_{i}$, so $v(G_{i})\geq v(G_{i-1})-2n_{i}/C\geq v(G_{i-1})-2n/C$. Thus, we have $v(G_{i})\geq n(1-2i/C)$ for all $i=1,\dots,I$. Furthermore, if the process did not halt at index $i$, we have $d_{i}\leq d_{i-1}/2$, and so $d_{i}\leq d2^{-i}$. This shows that $I\leq\log_{2}n$ and $n_{I}\geq n(1-2\log_{2}n/C)$.
Finally, note that $G_{I}$ has no vertices of degree more than $Cd(G_{I})$. Indeed, if the process has halted because $G_{I}$ contains no vertices of $Cd_{I}/2$, this is immediate, and if the process has halted because $d_{I}\geq d_{I-1}/2$, then $G_{I}$ contains no vertex of degree more than $Cd_{I-1}/2\leq Cd_{I}$. Hence, $G_{I}$ is a $C$-balanced induced subgraph of $G$ on at least $(1-2\log_{2}/C)n$ vertices. ∎
Now we are ready to prove the main result of this section. Conveniently, we do not have separate versions for the minimum eigenvalue and the MaxCut, as our result for the latter is already optimal.
Lemma 11.4.
Let $0<\varepsilon<\alpha$, then the following holds if $n$ is sufficiently large. Let $G$ be an $n$-vertex graph of edge density at least $1-n^{-\alpha}$. If $\operatorname{surp}^{*}(G)\leq n^{1+\varepsilon}$, then $G$ contains an induced subgraph of density at least $(1-O((\log n)^{2}n^{2\varepsilon-1}))$ on at least $n/2$ vertices.
@@
Let 1p1-p the edge density of $G$, and let $C=4\log_{2}n$. Applying Lemma 11.3 to the complement of $G$, we find an induced subgraph $G_{0}\subseteq G$ on at least $n/2$ vertices with edge density $1-p_{0}\geq 1-p$ such that the complement of $G_{0}$ is $C$-balanced. As $p_{0}\leq p\leq 0.001C^{-3}$, we can apply Lemma 11.2 to conclude that
$\operatorname{surp}^{*}(G_{0})=\Omega\left(\min\left\{\frac{n}{C^{3}p_{I}},C^{-1}p_{I}^{1/2}n^{3/2}\right\}\right).$
However, as $p_{0}<n^{-\alpha}$ and $\varepsilon<\alpha$, the inequality $\operatorname{surp}^{*}(G_{0})\leq n^{1+\varepsilon}$ is only possible if
$p_{0}\leq O(C^{2}n^{2\varepsilon-1})=O(n^{2\varepsilon-1}(\log n)^{2}).$
12 Finding large cliques
In this section, we combine our densification steps to prove that graphs with large smallest eigenvalue (or small surplus) contain large cliques.
@@ 12 Finding large cliques
Let $s\in(0,1/6)$ and $\delta>0$, then there exist $\rho>0$ such that the following holds for every sufficiently large $n$. Let $G$ be an $n$-vertex graph of edge density at least $n^{-\rho}$ such that $|\lambda_{n}|\leq n^{s}$. Then $G$ contains a clique of size at least $n^{1-2s-\delta}$.
Furthermore, we may choose $\rho=\alpha\delta$ for some decreasing $\alpha=\alpha(s)>0$.
@@
Let $\varepsilon=\min\{\delta/20,(1/6-s)/5\}$. Applying the phase 1 densification step, that is, Lemma 8.3, there exists $\rho=\rho(s,\varepsilon)>0$ and $c_{1}=c_{1}(s,\varepsilon)$ such that if $G$ has edge density at least $n^{-\rho}$ and the absolute value of the smallest eigenvalue $|\lambda_{n}|\leq n^{s}$, then $G$ contains a subgraph $G_{1}$ on at least $n_{1}\geq n^{1-\varepsilon}$ vertices with edge density at least $c_{1}$. Note that $G_{1}$ satisfies that if its smallest eigenvalue is $\lambda^{(1)}$, then $|\lambda^{(1)}|\leq|\lambda_{n}|\leq n^{1+s}\leq n_{1}^{(1+s)/(1-\varepsilon)}\leq n_{1}^{1+s_{1}}$ for $s_{1}=s+2\varepsilon\in(0,1/6)$. Furthermore, we may choose $\rho=\alpha_{0}\varepsilon$ for some $\alpha_{0}=\alpha_{0}(s)$, which shows that we may choose $\alpha=\alpha(s)>0$ such that $\rho=\alpha\delta$.
Next, applying the phase 2 densification to $G_{1}$, that is, Corollary 9.5, we get that $G_{1}$ contains a subgraph $G_{2}$ on $n_{2}\geq c_{1}n_{1}/2\geq n^{1-2\varepsilon}$ vertices with edge density at least $1-10^{-5}$. The smallest eigenvalue $\lambda^{(2)}$ of $G_{2}$ still satisfies that $|\lambda^{(2)}|\leq n_{2}^{1+s_{2}}$ for $s_{2}=s_{1}+\varepsilon=s+3\varepsilon\in(0,1/6)$.
Apply the phase 3 densification step to $G_{2}$, that is, Lemma 10.5. Then $G_{2}$ contains a subgraph $G_{3}$ of edge density $1-n_{2}^{3s_{2}-1+o(1)}>1-n_{2}^{-1/2}$ on $n_{3}\geq n_{2}^{1-o(1)}>n^{1-3\varepsilon}$ vertices. The smallest eigenvalue $\lambda^{(3)}$ of $G_{3}$ satisfies that $|\lambda^{(3)}|\leq n_{3}^{1+s_{3}}$ for $s_{3}=s_{2}+\varepsilon=s+4\varepsilon\in(0,1/6)$.
Finally, apply the phase 4 densification step to $G_{3}$, that is, Lemma 11.4. As $\operatorname{surp}^{*}(G_{3})\leq|\lambda^{(3)}|n_{3}\leq n_{3}^{1+s_{3}}$ and $s_{3}<1/6<1/2$, the lemma is indeed applicable. Therefore, we get a graph $G_{4}$ of density at least $1-O((\log n_{3})^{2}n_{3}^{2s_{3}-1})>1-n_{3}^{2s_{3}+\varepsilon-1}$ on at least $n_{3}/2>n^{1-4\varepsilon}$ vertices. But then by Turán’s theorem, $G_{4}$ contains a clique of size $n_{3}^{1-2s_{3}-\varepsilon}/4\geq n^{(1-4\varepsilon)(1-2s-9\varepsilon)}>n^{1-2s-13\varepsilon}>n^{1-2s-\delta}$, finishing the proof. ∎
Here, the bound $n^{1-2s-\delta}$ is optimal up to the $\delta$ error term. Indeed, the Erdős-Rényi random graph with edge probability $p=1-n^{2s-1}$ has no clique of size larger that $n^{1-2s+o(1)}$, and its smallest eigenvalue satisfies $|\lambda_{n}|=O(n^{s})$. Next, we use the previous theorem and Densification step 0 to prove Theorem 1.2, which we restate here.
Theorem 12.2.
There exists $c>0$ such that the following holds for every $s>0$, and every $d$ sufficiently large with respect to $s$. Let $G$ be a graph of average degree $d$ and assume that $|\lambda_{n}|\leq d^{s}$. Then $G$ contains a clique of size at least $d^{1-cs}$.
Proof.
Let $s_{0}=1/12$ and $\alpha=\alpha(s_{0})$ be the constant given by Theorem 12.1. Then $\alpha(s)\geq\alpha$ for every $s\leq s_{0}$. By densification step 0, that is Lemma 7.1, $G$ contains a subgraph $G_{0}$ on $d$ vertices of density at least $\Omega(1/|\lambda_{n}|)\geq\Omega(d^{-s})$. Set $\delta=2s/\alpha$, then the density of $G_{0}$ is at least $d^{-\alpha\delta}$, so we can apply Theorem 12.1 to $G_{0}$ to get a clique of size at least $d^{1-2s-\delta}=d^{1-s(2+2/\alpha)}$, so $c=\min\{12,2+2/\alpha\}$ suffices. ∎
Now we prove the MaxCut version of Theorem 12.1, and use it to deduce Theorem 1.3.
Theorem 12.3.
Let $s\in(0,1/60)$ and $\delta>0$, then there exist $\rho>0$ such that the following holds for every sufficiently large $n$. Let $G$ be an $n$-vertex graph of edge density at least $n^{-\rho}$ such that $\operatorname{surp}(G)\leq n^{1+s}$. Then $G$ contains a clique of size at least $n^{1-2s-\delta}$.
Proof.
@@
The proof of this is essentially identical to the proof of Theorem 12.1. The only difference is that we cite the MaxCut versions of our main densification results: Lemma 8.5, Corollary 9.7, Lemma 10.9 and Lemma 11.4. We also use that $\operatorname{surp}^{*}(G)\leq O(\operatorname{surp}(G)\log n)$ by Lemma 5.2. ∎
Proof of Theorem 1.3.
@@
Let $s=\delta_{0}=\min\{1/100,\delta/10\}$, and let $\rho$ be the constant guaranteed by Theorem 12.3 with $\delta_{0}$ instead of $\delta$. We show that $\varepsilon=\min\{\rho/3,s/2\}$ suffices. Let $G$ be a graph with $\operatorname{surp}(G)\leq m^{1/2+\varepsilon}$, and let $n$ be the number of vertices of $G$. We may assume that $G$ contains no isolated vertices. Then, a result of Erdős, Gyárfás, and Kohayakawa [36] implies that $\operatorname{surp}(G)\geq n/6$. If $m\leq n^{2-3\varepsilon}$, then $\operatorname{surp}(G)\geq n/6\geq\frac{1}{6}m^{1/(2-3\varepsilon)}\geq m^{1/2+\varepsilon}$, contradiction. Hence, $m\geq n^{2-3\varepsilon}$, so the edge density of $G$ is at least $n^{-3\varepsilon}>n^{-\rho}$. As $\operatorname{surp}(G)\leq m^{1/2+\varepsilon}\leq n^{1+2\varepsilon}\leq n^{1+s}$, $G$ contains a clique of size
$n^{1-2s-\delta_{0}}=n^{1-3s}\geq m^{(1-3s)/2}\geq m^{1/2-\delta}$
by Theorem 12.3. ∎

### v2

@@ 9.3 Further bounds on balanced graphs
First, assume that $N_{2}\leq\frac{1}{8}pn^{2}$. Note that $\mu_{1}^{2}+P_{2}+N_{2}=\|B\|_{F}^{2}$ is twice the number of edges of $\overline{G}$, so $\mu_{1}^{2}+P_{2}+N_{2}=2p\binom{n}{2}\geq pn^{2}/2$. Hence, using that $C^{2}p\leq 1/100$, we get
$P_{2}\geq pn^{2}/2-\mu_{1}^{2}-N_{2}\geq pn^{2}/2-C^{2}p^{2}n^{2}-pn^{2}/8\geq pn^{2}/4.$
But then, (12) implies $\operatorname{surp}^{*}(G)=\Omega(C^{-1/2}p^{1/2}n^{3/2})$, and we are done.
In the rest of the proof, we may assume $N_{2}\geq\frac{1}{8}pn^{2}$. By the inequality between the quadratic and cubic mean, we have
@@
Next, consider the quantity $T=N_{3}-P_{3}$. Observe that $\mu_{1}^{3}-T=\sum_{i=1}^{n}\mu_{i}^{3}$ is six-times the number of triangles of $\overline{G}$. In particular, $\mu_{1}^{3}-T$ it is nonnegative, showing that $T\leq\mu_{1}^{3}\leq\bar{\Delta}^{3}$. If $N_{3}\geq 2T$, or equivalently, $P_{3}\geq N_{3}/2$. We are done as (13) implies
$\operatorname{surp}^{*}(G)=\Omega\Big(\bar{\Delta}^{-1}{P_{3}}\Big)=\Omega\Big(\bar{\Delta}^{-1}{N_{3}}\Big)=\Omega\Big(C^{-1}p^{1/2}n^{3/2}\Big).$
@@
Finally, if $N_{3}\leq 2T$, then $\bar{\Delta}^{3}\geq T\geq N_{3}/2$. By the Cauchy–Schwartz inequality applied to the sequences $(|\mu_{i}|^{3})_{\mu_{i}<0}$ and $(|\mu_{i}|)_{\mu_{i}<0}$, we have $N_{1}N_{3}\geq N_{2}^{2}$, which gives
$N_{1}\geq\frac{N_{2}^{2}}{N_{3}}\geq\frac{(pn^{2}/8)^{2}}{2\bar{\Delta}^{3}}\geq\frac{n}{128C^{3}p}.$
But $0=\operatorname{tr}(B)=\mu_{1}+P_{1}-N_{1}$, from which
$P_{1}=N_{1}-\mu_{1}\geq\frac{n}{128C^{3}p}-\bar{\Delta}\geq\frac{n}{128C^{3}p}-Cpn\geq\frac{n}{256C^{3}p}.$
Here, we also used that $pC^{2}\leq 1/100$. This completes the proof by (11), i.e. $\operatorname{surp}^{*}(G)\geq P_{1}$. ∎
9.4 Completing the proof
We can summarize the results of Sections˜9.2 and 9.3 in the following theorem.
Theorem 9.9.
Let $G$ be an $n$-vertex graph with edge density 1p1-p, whose complement $\overline{G}$ is $C$-balanced of average degree $d=d(\overline{G})$. If $C^{2}p\leq 1/100$, then
  1. (a)
    the smallest eigenvalue $\lambda_{n}$ satisfies $|\lambda_{n}|=\Omega\Big(C^{-3}\cdot\max\big\{\min\{n/d,d^{1/2}\},d^{1/3}\big\}\Big),$
  2. (b)
    and the surplus satisfies $\operatorname{surp}^{*}(G)=\Omega\Big(C^{-3}n\cdot\max\big\{\min\{n/d,d^{1/2}\},d^{1/4}\big\}\Big).$
Combining Lemma˜9.2 and Theorem˜9.9, we now prove the main theorem of this section.
Proof of Theorem 9.1.
(a) Applying Lemma˜9.2 to $\overline{G}$, we get that $G$ contains an induced subgraph $G_{1}$ on $n_{1}=\Omega(n/\log n)$ vertices of density $1-p_{1}\geq 1-10^{-6}$, and whose complement $\overline{G_{1}}$ is $C_{1}$-balanced for $C_{1}=4\log_{2}1/p_{1}$. We are done if $G_{1}$ is a clique. Otherwise, $p_{1}\geq 1/\binom{n_{1}}{2}$ implying that $C_{1}=O(\log n_{1})$. We claim that $G_{1}$ is the desired subgraph, i.e. $p_{1}=\widetilde{O}\big(|\lambda_{n}|^{2}/n\big)$, by applying Theorem˜9.9 (note that this is valid as as $C_{1}^{2}p_{1}\leq 16p_{1}(\log_{2}1/p_{1})^{2}\leq 1/100$ for all $p_{1}\in(0,10^{-6})$).
Let $\lambda_{\min}$ be the smallest eigenvalue of $G_{1}$. Cauchy’s interlacing theorem guarantees $|\lambda_{\min}|\leq|\lambda_{n}|\leq n^{1/4}/(\log n)^{4}$. Let $d_{1}=d(\overline{G_{1}})$. Suppose $d_{1}\geq n_{1}^{2/3}$. It is easy to check that $\max\big\{\min\{n_{1}/d_{1},d_{1}^{1/2}\},d_{1}^{1/3}\big\}\geq n_{1}^{1/4}$. Indeed, $d_{1}^{1/3}\geq n_{1}^{1/4}$ if $d_{1}\geq n_{1}^{3/4}$ and $\min\{n_{1}/d_{1},d_{1}^{1/2}\}\geq n_{1}^{1/4}$ if $n_{1}^{2/3}\leq d_{1}\leq n_{1}^{3/4}$. So Theorem˜9.9 implies $|\lambda_{\min}|=\Omega\big(C_{1}^{-3}\cdot n_{1}^{1/4}\,\big)$. Recalling that $n_{1}=\Omega(n/\log n),C_{1}=O(\log n_{1})$ and $n$ is sufficiently large, this bound gives $|\lambda_{\min}|=\Omega(n^{1/4}/(\log n)^{13/4}\,\big)>n^{1/4}/(\log n)^{4}$; this is impossible.
Now we know that $d_{1}\leq n_{1}^{2/3}$. In this case, $d_{1}^{1/3}\leq d_{1}^{1/2}\leq n_{1}/d_{1}$, so Theorem˜9.9 implies $|\lambda_{\min}|=\Omega\big(C_{1}^{-3}\cdot d_{1}^{1/2}\,\big)$. Recall that $|\lambda_{\min}|\leq|\lambda_{n}|$. It must be that $C_{1}^{-3}\cdot d_{1}^{1/2}=O\big(|\lambda_{n}|\big)$, i.e. $d_{1}=O\big(|\lambda_{n}|^{2}C_{1}^{6}\,\big)=O\big(|\lambda_{n}|^{2}(\log n)^{6}\big)$. This completes the proof as $p_{1}=d_{1}/(n_{1}-1)=O\big(|\lambda_{n}|^{2}(\log n)^{7}/n\big)$.
(b) By Lemma˜9.2 on $\overline{G}$, $G$ has an induced subgraph $G_{1}\subset G$ on $n_{1}=\Omega(n/\log n)$ vertices of density $1-p_{1}\geq 1-10^{-6}$ and with a $C_{1}$-balanced complement, where $C_{1}=O(\log n)$. By Lemma˜5.2, $\operatorname{surp}^{*}(G_{1})\leq\operatorname{surp}^{*}(G)=O\big(\operatorname{surp}(G)\cdot\log n\big)=O(n^{6/5}/(\log n)^{5})$. Write $d_{1}=d(\overline{G_{1}})$. If $d_{1}\geq n_{1}^{2/3}$, one can check that $\max\big\{\min\{n_{1}/d_{1},d_{1}^{1/2}\},d_{1}^{1/4}\big\}\geq n_{1}^{1/5}$, so Theorem˜9.9 yields $\operatorname{surp}^{*}(G_{1})=\Omega\big(C_{1}^{-3}n_{1}\cdot n_{1}^{1/5}\big)=\Omega\big(n_{1}^{6/5}/C_{1}^{3}\big)=\Omega\big(n^{6/5}/(\log n)^{6/5+3}\big)>n^{6/5}/(\log n)^{5}$; this is impossible. So $d_{1}\leq n_{1}^{2/3}$ must hold. But then, Theorem˜9.9 implies $\operatorname{surp}^{*}(G_{1})=\Omega\big(C_{1}^{-3}n_{1}\cdot d_{1}^{1/2}\big)$. As $\operatorname{surp}^{*}(G_{1})\leq\operatorname{surp}^{*}(G)=O(\operatorname{surp}(G)\cdot\log n)$, we get $d_{1}=O\big(C_{1}^{6}\operatorname{surp}(G)^{2}(\log n)^{2}/n_{1}^{2}\big)$. So, $p_{1}=d_{1}/(n_{1}-1)=\widetilde{O}(\operatorname{surp}(G)^{2}/n^{3})$, as desired. ∎
10 Large cliques from small eigenvalues or surplus
In this section, we combine our densification steps to prove that graphs with large smallest eigenvalue (or with small surplus) contain large cliques. This section contains two main results: one result covers graphs whose density is polynomially close to $n$, while the other covers graphs whose average degree could be much smaller than $n$.
Theorem 10.1.
Let $\gamma,\varepsilon,\rho>0$ such that $\rho<1/2$, $\varepsilon+6\gamma<1$ and $\rho/\varepsilon+2\gamma/(1-\varepsilon-4\gamma)<1$. Then for sufficiently large $n$, any $n$-vertex graph $G$ of edge density at least $n^{-\rho}$ and smallest eigenvalue $|\lambda_{n}|\leq n^{\gamma}$ contains a clique of size at least $n^{1-\varepsilon-2\gamma}$.
@@ 10 Large cliques from small eigenvalues or surplus
Choose a parameter $\varepsilon^{\prime}<\varepsilon$ for which the inequalities
$\rho<1/2,\quad\varepsilon^{\prime}+6\gamma<1\text{\ \ and\ \ }\rho/\varepsilon^{\prime}+2\gamma/(1-\varepsilon^{\prime}-4\gamma)<1$
are still satisfied. This can be done since all the inequalities are strict.
As the first step, we apply Lemma˜7.3, which states that if the above inequalities on the parameters $\gamma,\varepsilon^{\prime}$ and $\rho$ are satisfied and $n$ is sufficiently large, there exists an induced subgraph $G_{1}\subseteq G$ on at least $n_{1}\geq n^{1-\varepsilon^{\prime}}$ vertices with density $\Omega(1)$ (where the hidden constant may depend on $\gamma,\rho$ and $\varepsilon^{\prime}$). Denote the smallest eigenvalue of $G_{1}$ by $\lambda^{(1)}$; then Cauchy’s interlacing theorem implies $|\lambda^{(1)}|\leq|\lambda_{n}|\leq n^{\gamma}$.
Next, applying the second phase of densification to $G_{1}$, that is, Corollary˜8.6, we get that $G_{1}$ contains an induced subgraph $G_{2}$ on $n_{2}=\Omega(n_{1})=\Omega(n^{1-\varepsilon^{\prime}})$ vertices with edge density at least $1-10^{-6}$. The smallest eigenvalue $\lambda^{(2)}$ of $G_{2}$ still satisfies $|\lambda^{(2)}|\leq|\lambda^{(1)}|\leq n^{\gamma}$.
Finally, we apply the third phase of densification to $G_{2}$, that is, Theorem˜9.1. This theorem states that if $|\lambda^{(2)}|\leq n_{2}^{1/4}/(\log n)^{4}$, then $G_{2}$ contains an induced subgraph $G_{3}$ on $n_{3}=\Omega(n_{2}/\log n_{2})$ vertices with edge density $1-p_{3}$, where $p_{3}\leq|\lambda^{(2)}|^{2}/n_{2}^{1-o(1)}$. The condition on $\lambda^{(2)}$ is easy to verify since $|\lambda^{(2)}|\leq n^{\gamma}=o\big(n^{(1-\varepsilon^{\prime})/4}/(\log n)^{4}\big)<n_{2}^{1/4}/(\log n_{2})^{4}$. Here, we used $n_{2}=\Omega(n^{1-\varepsilon^{\prime}})$ and $4\gamma<1-\varepsilon^{\prime}$, which follows directly from $\varepsilon^{\prime}+6\gamma<1$.
To conclude the proof, we apply Turán’s theorem to $G_{3}$. This guarantees $G_{3}$ has a clique of size
$\min\Big(n_{3},\Omega\big(p_{3}^{-1}\big)\Big)\geq\frac{n_{2}^{1-o(1)}}{|\lambda^{(2)}|^{2}}\geq\frac{n^{1-\varepsilon^{\prime}-o(1)}}{n^{2\gamma}}\geq n^{1-\varepsilon^{\prime}-2\gamma-o(1)}.$
Since $\varepsilon^{\prime}<\varepsilon$, we conclude that $G$ contains a clique of size at least $n^{1-\varepsilon-2\gamma}$, if $n$ is sufficiently large. ∎
Here, the bound $n^{1-2\gamma-\varepsilon}$ is close to optimal (when $\rho,\varepsilon$ are both small). Indeed, the Erdős-Rényi random graph $G(n,p)$ with edge probability $p=1-n^{2\gamma-1}$ has no clique of size larger than $n^{1-2\gamma+o(1)}$, and its smallest eigenvalue satisfies $|\lambda_{n}|=\Theta(n^{\gamma})$ with high probability. We now shift our focus to sparse graphs, with the goal of proving Theorem 1.2, which we restate here.
Theorem 10.2.
Let $\gamma\in(0,1/10)$, and let $d$ be sufficiently large. Then, every graph $G$ of average degree $d$ whose smallest eigenvalue satisfies $|\lambda_{n}|\leq d^{\gamma}$ contains a clique of size at least $d^{1-4\gamma}$.
@@
We will reduce to Theorem˜10.1 by finding an induced subgraph $G_{0}\subseteq G$ whose average degree is polynomially close to the number of its vertices. To do this, let $x$ be a vertex of maximum degree in $G$, and let $S$ be a set of $d$ neighbours of $x$. We claim that $e(G[S])\geq d^{2}/4|\lambda_{n}|$, and then will take $G_{0}=G[S]$.
To prove the claim, define a vector $v\in\mathbb{R}^{n}$ such that $v(x)=1$, $v(y)=\frac{\lambda_{n}}{d}$ for $y\in S$, and $v(z)=0$ if $z\not\in S\cup\{x\}$. Since $\lambda_{n}$ is the smallest eigenvalue of the adjacency matrix $A$, we must have $v^{T}Av\geq\lambda_{n}\|v\|_{2}^{2}$. For this $v$, we also have $v^{T}Av=2\lambda_{n}+2\frac{\lambda_{n}^{2}}{d^{2}}e(G[S])$ and $\|v\|_{2}^{2}=1+d\frac{\lambda_{n}^{2}}{d^{2}}\leq 3/2$. Combining these two observations, we have
$2\lambda_{n}+2\frac{\lambda_{n}^{2}}{d^{2}}e(G[S])=v^{T}Av\geq\lambda_{n}\|v\|_{2}^{2}\geq\frac{3}{2}\lambda_{n}.$
Therefore, we conclude that $e(G[S])\geq\frac{d^{2}}{4|\lambda_{n}|}$, as needed.
Now, define $G_{0}=G[S]$; it has $d$ vertices and density at least $d^{-\gamma}/2$. Hence, Theorem˜10.1 applies to $G_{0}$ with parameters $\rho=\gamma+\kappa$ and $\varepsilon=2\gamma$, where $\kappa>0$ is an arbitrary small constant, assuming $d$ is sufficiently large. These parameters clearly satisfy $\rho<1/2$ and $\varepsilon+6\gamma=8\gamma<1$ (using $\gamma<1/10$) and we also have $\rho/\varepsilon+2\gamma/(1-\varepsilon-4\gamma)=1/2+2\gamma/(1-6\gamma)+\kappa/\varepsilon<1$ for $\kappa$ sufficiently small with respect to $\gamma$. Hence, we conclude that $G_{0}$ contains a clique of size $d^{1-\varepsilon-2\gamma}=d^{1-4\gamma}$. This completes the proof. ∎
Now we prove the MaxCut version of Theorem˜10.1, and use it to deduce Theorem˜1.3.
@@
Let $\gamma\in(0,1/60)$ and $\delta>0$, then there exist $\rho>0$ such that the following holds for every sufficiently large $n$. Let $G$ be an $n$-vertex graph of edge density at least $n^{-\rho}$ such that $\operatorname{surp}(G)\leq n^{1+\gamma}$. Then $G$ contains a clique of size at least $n^{1-2\gamma-\delta}$.
@@
The proof of this is essentially identical to the proof of Theorem˜10.1. The only difference is that we cite the MaxCut versions of our main densification results: Lemma˜7.5, Corollary˜8.8, and (b) of Theorem˜9.1. ∎
Proof of Theorem˜1.3.
@@
Let $\gamma=\delta/2$, and let $\rho$ be the constant in Theorem˜10.3. We show that $\varepsilon=\min\{\rho/4,\gamma/2\}$ suffices. Suppose for contradiction that $\operatorname{surp}(G)<m^{0.5+\varepsilon}$. Let $n$ be the number of vertices in $G$. We may assume that $G$ contains no isolated vertices. Then, a result of Erdős, Gyárfás, and Kohayakawa [41] shows that $\operatorname{surp}(G)\geq\frac{n}{6}$, so $m>(n/6)^{1/(0.5+\varepsilon)}>n^{2-4\varepsilon}$. Then edge density of $G$ is at least $n^{-4\varepsilon}\geq n^{-\rho}$. As $m\leq n^{2}$, $\operatorname{surp}(G)\leq m^{0.5+\varepsilon}\leq n^{1+2\varepsilon}\leq n^{1+\gamma}$. We get the contradiction as Theorem˜10.3 guarantees that $G$ contains a clique of size $n^{1-2\gamma-\delta}=n^{1-2\delta}\geq m^{1/2-\delta}$. ∎